# Armijo Line-search Makes (Stochastic) Gradient Descent Go Fast

### v1

@@ Abstract
Armijo line-search (Armijo-LS) is a standard method to set the step-size for gradient descent (GD). For smooth functions, Armijo-LS alleviates the need to know the global smoothness constant $L$ and adapts to the “local” smoothness, enabling GD to converge faster. However, existing theoretical analyses of GD with Armijo-LS (GD-LS) do not characterize this fast convergence. We show that if the objective function satisfies a certain non-uniform smoothness condition, GD-LS converges provably faster than GD with a constant $1/L$ step-size (denoted as GD(1/L)). Our results imply that for convex losses corresponding to logistic regression and multi-class classification, GD-LS can converge to the optimum at a linear rate, and hence, improve over the sublinear convergence of GD(1/L). Furthermore, for non-convex losses satisfying gradient domination (for example, those corresponding to the softmax policy gradient in RL or generalized linear models with a logistic link function), GD-LS can match the fast convergence of algorithms tailored for these specific settings. Finally, we prove that under the interpolation assumption, for convex losses, stochastic GD with a stochastic line-search can match the fast convergence of GD-LS.
@@ 1 Introduction
One often-cited reason to explain the faster convergence of GD-LS is that it adapts to the “local” smoothness constant $L(\theta)$ near the point $\theta$, and results in an effective step-size of $\nicefrac{{1}}{{L(\theta)}}$. In some regions, $L(\theta)$ might be much smaller than the global smoothness $L$, thus allowing GD-LS to use much bigger step-sizes and consequently lead to faster convergence. However, existing theoretical analyses of GD-LS do not formalize this intuition, and can therefore, not explain the algorithm’s faster convergence.
In this paper, we aim to solve $\min_{\theta\in\mathbb{R}^{d}}f(\theta)$ for a special class of objective functions and formally characterize the advantage of GD-LS over GD(1/L). In particular, we make the following contributions.
@@
Contribution 1. In Section 2, we introduce a class of functions that satisfy an $(L_{0},L_{1})$ non-uniform smoothness condition. This condition is similar to that proposed in Zhang et al. (2019) to explain the success of normalization and gradient clipping techniques when training neural networks. In particular, we consider functions where the local smoothness constant around a point $\theta$ is given by $L(\theta)=L_{0}+L_{1}\,f(\theta)$ where $L_{0}$ and $L_{1}$ are non-negative constants ($L_{1}=0$ corresponds to the standard uniform smoothness). We show that the proposed condition is satisfied by common objectives; for example, both the logistic and exponential losses used for linear classification satisfy the condition with $L_{0}=0$ and $L_{1}\neq 0$. Furthermore, we prove that this condition is also satisfied by non-convex functions corresponding to generalized linear models with a logistic link function and the softmax policy gradient objective in reinforcement learning.
@@
Contribution 2. In Section 3, we analyze the convergence of GD-LS on functions satisfying the proposed conditions. For this, we first prove that the step-size selected by Armijo-LS around $\theta$ is lower-bounded by $1/L(\theta)$, and hence, GD-LS provably adapts to the local smoothness. We use this property to prove Theorem 1, a meta-theorem that quantifies the convergence rate of GD-LS. We note that Hübler et al. (2024) also consider GD-LS for minimizing a different class of non-uniform smooth functions. However, their analysis for general non-convex functions does not demonstrate the algorithm’s adaptivity to the smoothness, nor does it result in a faster rate than GD(1/L). Moreover, their resulting algorithm requires the knowledge of the non-uniform smoothness constant, making it impractical.
@@
Contribution 3. In Section 4, we instantiate Theorem 1 for non-uniform smooth, convex losses that include logistic regression and multi-class classification with the cross-entropy loss as examples. Specifically, in Corollary 1, we show that GD-LS converges at an $O\left(\left(\nicefrac{{f^{*}}}{{\epsilon}}\right)\,\ln\left(\nicefrac{{1}}{{\epsilon}}\right)\right)$ rate where $f^{*}:=\inf f(\theta)$. Hence, when $f^{*}$ is $O(\epsilon)$, GD-LS converges at an $O\left(\ln\left(\nicefrac{{1}}{{\epsilon}}\right)\right)$ rate, compared to the sublinear $O\left(\nicefrac{{1}}{{\epsilon}}\right)$ convergence of GD(1/L). We instantiate this result for logistic regression on linearly separable data, and prove the linear convergence of GD-LS (Corollary 7), thus matching the rate for normalized GD (Axiotis & Sviridenko, 2023).
@@
Contribution 4. In Section 5, we instantiate Theorem 1 for non-convex functions satisfying non-uniform smoothness and gradient domination conditions that guarantee global optimality. Specifically, in Section 5.1, we analyze the convergence of GD-LS on the softmax policy gradient objective in reinforcement learning (Mei et al., 2020). In this setting, the linear convergence rate attained by GD-LS is provably better than the $\Omega(\nicefrac{{1}}{{\epsilon}})$ convergence of GD(1/L) and matches the rate of natural policy gradient (Kakade & Langford, 2002). In Section 5.2, we analyze the convergence of GD-LS for functions satisfying the PL condition (Polyak, 1987; Karimi et al., 2016) and instantiate the result for generalized linear models with the logistic link function. Our result demonstrates that GD-LS can converge faster than both GD and normalized GD (Hazan et al., 2015).
@@
Contribution 5. Finally, in Section 6, we show that the advantages of line-search carry over to the stochastic setting. Specifically, we consider a finite-sum objective $f(\theta)=\frac{1}{n}\sum_{i=1}^{n}\,f_{i}(\theta)$, and study the convergence of stochastic gradient descent (SGD) in conjunction with a stochastic line-search (Vaswani et al., 2019b). We restrict our attention to the interpolation setting (Vaswani et al., 2019a; Ma et al., 2018; Schmidt & Roux, 2013) which implies that each $f_{i}$ is minimized at $\theta^{*}:=\operatorname*{arg\,min}f(\theta)$. Such a condition is satisfied by logistic regression on linearly separable data. Interpolation enables the fast convergence of SGD, allowing it to match the GD rate but with an $O(1)$ iteration cost. Under interpolation, SGD with a stochastic line-search (referred to as SGD-SLS) and its variants empirically outperform constant step-size SGD, and have been used to train deep neural networks (Vaswani et al., 2019b; Galli et al., 2024). We provide further theoretical justification for the fast convergence of SGD-SLS. Specifically, in  Corollary 5, we prove that for logistic regression on linearly separable data, SGD-SLS converges at a linear rate.
@@ 2 Problem Formulation
We aim to solve the unconstrained minimization problem: $\min_{\theta\in\mathbb{R}^{d}}f(\theta)$. We define $\theta^{*}\in\arg\inf f(\theta)$ as an optimal solution and $f^{*}:=\inf f(\theta)$ as the minimum function value. Throughout, we consider $f$ to be twice-differentiable and satisfies the following assumptions:
@@
$f$ is $(L_{0},L_{1})$ non-uniform smooth i.e. for $L_{0},L_{1}\geq 0$,
(a) For all $x,y$ such that $\left\|x-y\right\|\leq\frac{q}{L_{1}}$ where $q\geq 1$ is a constant, if $A:=1+e^{q}-\frac{e^{q}-1}{q}$ and $B:=\frac{e^{q}-1}{q}$,
@@
If $L_{1}=0$Assumption 2 recovers the standard uniform smoothness condition as a special case. Consequently, common smooth objectives such as linear regression or logistic regression satisfy the above condition. For example, if $X\in\mathbb{R}^{d\times n}$ is the feature matrix, and $y\in\mathbb{R}^{n}$ is the vector of measurements, then the linear regression objective $f(\theta)=\frac{1}{2n}\,\left\|X\,\theta-y\right\|_{2}^{2}$ is $(\frac{1}{n}\lambda_{\max}[X^{T}X],0)$ non-uniform smooth where $\lambda_{\max}[A]$ is the maximum eigenvalue of the PSD matrix $A$. The above condition is similar to the non-uniform smoothness conditions proposed in the literature (Zhang et al., 2019, 2020; Chen et al., 2023). The difference is that the smoothness in Assumption 2 is proportional to $f(\theta)$ rather than $\left\|\nabla f(\theta)\right\|$ in the previous work. Subsequently, we will see that GD with Armijo line-search is easier to analyze with this alternate definition of non-uniform smoothness.
In order to show the benefit of GD with Armijo line-search, we focus on functions where $L_{1}\neq 0$. We will require these functions to satisfy an additional assumption that relates the gradient norm to the function value. As we will see, such an assumption is typically true for losses with an exponential tail, even when using a 2 layer neural network (Taheri & Thrampoulidis, 2023; Wu et al., 2024).
@@
For all $\theta$, there exist constants $\omega,\nu>0$ s.t.
@@
To motivate the above assumptions, we prove that common convex objectives for supervised learning such as linear logistic regression and linear multi-class classification satisfy the above condition with $L_{0}=0$ and non-zero $L_{1}$. Moreover, we show that these functions also satisfy Assumption 3 (all proofs are deferred to Appendix A). Below, we state the result for logistic regression and defer the results for the other losses to the Appendix.
@@
satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$, and Assumption 3 with $\nu=\max_{i}\left\|x_{i}\right\|$ and $\omega=0$.
Note that the logistic regression objective is also uniform smooth, meaning that it simultaneously satisfies Assumption 2 with $L_{0}=\frac{1}{4n}\,\lambda_{\max}[X^{T}X]$ and $L_{1}=0$, where $X\in\mathbb{R}^{n\times d}$ is the corresponding feature matrix. On the other hand, binary classification with the exponential loss is not uniform smooth on an unbounded domain, but satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$ (see Proposition 5 in Appendix A). We note that for logistic regression, the loss corresponding to a single point $i\in[n]$ satisfies the notion of non-uniform smoothness in Zhang et al. (2019) with $L_{0}=0$ and $L_{1}=\left\|x_{i}\right\|$ (Gorbunov et al., 2024, Example 1.6). However, $f(\theta)$ does not necessarily satisfy this assumption with $L_{0}=0$ (see Proposition 7 for a simple counter-example). Subsequently, we will see that having $L_{0}=0$ is key to achieving fast convergence for GD-LS. This further motivates our alternate definition of non-uniform smoothness.
Next, we show that the non-convex objective corresponding to generalized linear models (GLM) with a logistic link function also satisfies Assumptions 1, 2 and 3 . In particular, we prove the following proposition in Appendix A.
@@
Consider $n$ points where $x_{i}\in\mathbb{R}^{d}$ are the features and $y_{i}\in[0,1]$ are the corresponding labels. If $\pi_{i}(\theta)=\sigma(\langle x_{i},\theta\rangle):=\frac{1}{1+\exp(-\langle x_{i},\theta\rangle)}$, the GLM objective
$f(\theta)=\frac{1}{2n}\sum_{i=1}^{n}\left(\pi_{i}(\theta)-y_{i}\right)^{2}$ (4)
satisfies Assumption 2 with $L_{0}=\frac{17}{16}\,\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$ and $L_{1}=2\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$ and Assumption 3 with $\nu=2\max_{i}\left\|x_{i}\right\|$ and $\omega=\max_{i}\left\|x_{i}\right\|$.
Finally, in Appendix A, we also show that the objective for softmax policy gradient (Mei et al., 2020) also satisfies the required assumptions for multi-armed bandits and tabular Markov decision processes, and study these settings in more detail in Section 5.
Now that we have motivated the use of consider Assumptions 1 and 3, in the next section, we will consider minimizing non-uniform smooth functions using gradient descent.

### v2

# Armijo Line-search Can Make (Stochastic) Gradient Descent Provably Faster

@@ Abstract
Armijo line-search (Armijo-LS) is a standard method to set the step-size for gradient descent (GD). For smooth functions, Armijo-LS alleviates the need to know the global smoothness constant $L$ and adapts to the “local” smoothness, enabling GD to converge faster. Existing theoretical analyses show that GD with Armijo-LS (GD-LS) can result in constant factor improvements over GD with a $1/L$ step-size (denoted as GD(1/L)). We strengthen these results and show that if the objective function satisfies a certain non-uniform smoothness condition, GD-LS can result in a faster convergence rate than GD(1/L). In particular, we prove that for convex objectives corresponding to logistic regression and multi-class classification, GD-LS can converge to the optimum at a linear rate, and hence improves over the sublinear convergence of GD(1/L). Furthermore, for non-convex objectives satisfying gradient domination (e.g. those corresponding to the softmax policy gradient in RL or generalized linear models with a logistic link function), GD-LS can match the fast convergence of algorithms tailored for these specific settings. Finally, we prove that under the interpolation assumption, for convex losses, stochastic GD with a stochastic line-search can match the fast convergence of GD-LS.
@@ 1 Introduction
One often-cited reason to explain the faster convergence of GD-LS is that it adapts to the “local” smoothness constant $L(\theta)$ near the point $\theta$, and results in an effective step-size of $\nicefrac{{1}}{{L(\theta)}}$. In some regions, $L(\theta)$ might be much smaller than the global smoothness $L$, thus allowing GD-LS to use much bigger step-sizes and consequently lead to faster convergence. Existing works that define notions of local smoothness (Scheinberg et al., 2014; Lu & Mei, 2023; Mishkin et al., 2024; Fox & Schmidt, , and references there in) can be used to formalize this intuition and show that GD-LS can result in constant factor improvements over GD(1/L), while having the same rate of convergence. In contrast, we consider a special class of objective functions and show that GD-LS (with a large initial step-size) can result in a provably faster rate of convergence compared to GD(1/L). In particular, we make the following contributions.
@@
Contribution 1. In Sec. 2, we introduce a class of functions that satisfy an $(L_{0},L_{1})$ non-uniform smoothness condition. In particular, we consider functions where the local smoothness constant around a point $\theta$ is given by $L(\theta)=L_{0}+L_{1}\,f(\theta)$ where $L_{0}$ and $L_{1}$ are non-negative constants (and $L_{1}=0$ corresponds to the standard uniform smoothness). We show that the proposed condition is satisfied by common objectives; for example, both the logistic and exponential losses used for linear classification satisfy the condition with $L_{0}=0$ and $L_{1}\neq 0$. Furthermore, we prove that this condition is also satisfied by non-convex functions corresponding to generalized linear models with a logistic link function, and the softmax policy gradient objective in reinforcement learning.
This condition is similar to that proposed in Zhang et al. (2019) to explain the success of normalization and gradient clipping when training neural networks. The difference is that we consider problems where the smoothness is proportional to $f(\theta)$ rather than $\left\|\nabla f(\theta)\right\|$ (Zhang et al., 2019, 2020; Chen et al., 2023). Furthermore, in Sec. 2, we show that non-negative, twice-differentiable functions satisfying the non-uniform smoothness condition in Zhang et al. (2019) satisfy the proposed condition.
@@
Contribution 2. In Sec. 3, we analyze the convergence of GD-LS on functions satisfying the proposed conditions. We first prove that the step-size selected by GD-LS around $\theta$ is lower-bounded by $1/L(\theta)$, and hence, GD-LS provably adapts to the local smoothness. We use this to prove Thm. 1, a meta-theorem that quantifies the convergence rate of GD-LS for both convex and non-convex objectives.
@@
Contribution 3. In Sec. 4, we instantiate Thm. 1 for non-uniform smooth, convex losses that include logistic regression and multi-class classification with the cross-entropy loss as examples. Specifically, in Cor. 1, we show that GD-LS (with a sufficiently large initial step-size) converges at an $O\left(\left(\nicefrac{{f^{*}}}{{\epsilon}}\right)\,\ln\left(\nicefrac{{1}}{{\epsilon}}\right)\right)$ rate where $f^{*}:=\inf f(\theta)$. Hence, when $f^{*}$ is $\Theta(\epsilon)$, GD-LS converges at an $O\left(\ln\left(\nicefrac{{1}}{{\epsilon}}\right)\right)$ rate, compared to the sublinear $O\left(\nicefrac{{1}}{{\epsilon}}\right)$ convergence of GD(1/L). We further instantiate this result for logistic regression on linearly separable data, and prove the linear convergence of GD-LS (Cor. 2), thus matching the rate for normalized GD (Axiotis & Sviridenko, 2023). In App. C, we show that GD with the Polyak step-size (Polyak, 1987) can inherit this fast convergence for logistic regression.
In comparison to our results, most work exploiting non-uniform smoothness (Zhang et al., 2019, 2020; Koloskova et al., 2023; Chen et al., 2023; Li et al., 2023) requires the knowledge of the corresponding problem-dependent constants. Notable exceptions include Hübler et al. (2024) and (Vankov et al., 2024; Takezawa et al., 2024; Gorbunov et al., 2024) that consider GD-LS and the Polyak step-size respectively, and aim to minimize the class of non-uniform smooth functions in (Zhang et al., 2019). In particular, the work in Hübler et al. (2024) considers general non-convex functions and does not demonstrate the algorithm’s adaptivity to the smoothness, nor does it result in a faster rate than GD(1/L). Moreover, their resulting algorithm requires the knowledge of the non-uniform smoothness constant, making it impractical. On the other hand, (Vankov et al., 2024; Takezawa et al., 2024; Gorbunov et al., 2024) consider minimizing convex, non-uniform smooth functions using GD with the Polyak step-size. In Sec. 4, we show that GD-LS (without any modification) can achieve a similar convergence guarantee as these works.
@@
Contribution 4. In Sec. 5, we instantiate Thm. 1 for non-convex functions satisfying non-uniform smoothness and gradient domination conditions that guarantee global optimality. Specifically, in Sec. 5.1, we analyze the convergence of GD-LS for the softmax policy gradient objective in reinforcement learning (Mei et al., 2020). In this setting, the linear convergence rate attained by GD-LS is provably better than the $\Omega(\nicefrac{{1}}{{\epsilon}})$ convergence of GD(1/L) and matches the rate of natural policy gradient (Kakade & Langford, 2002). In Sec. 5.2, we analyze the convergence of GD-LS for functions satisfying the PL condition (Polyak, 1987; Karimi et al., 2016) and instantiate the result for generalized linear models with the logistic link function. Our result demonstrates that GD-LS can converge faster than both constant step-size and normalized GD (Hazan et al., 2015).
@@
Contribution 5. Finally, in Sec. 6, we show that the advantages of line-search carry over to the stochastic setting. Specifically, we consider a finite-sum objective $f(\theta)=\frac{1}{n}\sum_{i=1}^{n}\,f_{i}(\theta)$, and study the convergence of stochastic gradient descent (SGD) in conjunction with a stochastic line-search proposed in Vaswani et al. (2019b). We restrict our attention to the interpolation setting (Vaswani et al., 2019a; Ma et al., 2018; Schmidt & Roux, 2013) which implies that each $f_{i}$ is minimized at $\theta^{*}:=\operatorname*{arg\,min}f(\theta)$. The interpolation assumption is of practical interest (Zhang et al., 2017; Belkin et al., 2019), for example, it is satisfied for logistic regression with linearly separable data.
Interpolation enables the fast convergence of SGD, allowing it to match the GD convergence rate, but with an $O(1)$ iteration cost. Under interpolation, SGD with a stochastic line-search (referred to as SGD-SLS) and its variants empirically outperform constant step-size SGD, and have been used to train deep neural networks (Vaswani et al., 2019b; Galli et al., 2024). We provide further theoretical justification for the fast convergence of SGD-SLS. Specifically, in Cor. 5, we prove that for logistic regression with linearly separable data, SGD-SLS can converge at a linear rate.
@@ 2 Problem Formulation
We aim to solve the unconstrained minimization problem: $\min_{\theta\in\mathbb{R}^{d}}f(\theta)$. We define $\theta^{*}\in\arg\inf f(\theta)$ as an optimal solution and $f^{*}:=\inf f(\theta)$ as the minimum function value. Throughout, we consider $f$ to be twice-differentiable and satisfying the following assumptions:
@@
$f$ is $(L_{0},L_{1})$ non-uniform smooth i.e. for constants $L_{0},L_{1}\geq 0$,
(a) For all $x,y$ such that $\left\|x-y\right\|\leq\frac{q}{L_{1}}$, where $q\geq 1$ is a constant, if $A:=1+e^{q}-\frac{e^{q}-1}{q}$ and $B:=\frac{e^{q}-1}{q}$,
@@
The above condition is similar to the non-uniform smoothness conditions proposed in the literature (Zhang et al., 2019, 2020; Chen et al., 2023). Subsequently, we will see that this alternate definition of non-uniform smoothness is more general and enables a simpler analysis for GD-LS.
If $L_{1}=0$Assn. 2 recovers the standard uniform smoothness condition as a special case. Consequently, common smooth objectives such as linear regression or logistic regression satisfy the above condition. For example, if $X\in\mathbb{R}^{d\times n}$ is the feature matrix, and $y\in\mathbb{R}^{n}$ is the vector of measurements, then the linear regression objective, $f(\theta)=\frac{1}{2n}\,\left\|X\,\theta-y\right\|_{2}^{2}$ is $(\frac{1}{n}\lambda_{\max}[X^{T}X],0)$ non-uniform smooth where $\lambda_{\max}[A]$ is the maximum eigenvalue of the positive semi-definite matrix $A$.
In order to show the benefit of GD with Armijo line-search, we will focus on functions where $L_{1}\neq 0$, and the smoothness depends on the function value. We will require these functions to satisfy an additional assumption that relates the gradient norm to the function value. As we will see, such an assumption is true for losses with an exponential tail, even when using a 2 layer neural network (Taheri & Thrampoulidis, 2023; Wu et al., 2024).
@@
For all $\theta$, there exist constants $\omega,\nu\geq 0$ s.t.
@@
To motivate the above assumptions, we prove that common convex objectives for supervised learning such as linear logistic regression and linear multi-class classification satisfy Assn. 2 with $L_{0}=0$ and non-zero $L_{1}$. Moreover, we show that these functions also satisfy Assn. 3 with $\omega=0$. Below, we state the result for logistic regression, and defer the other results and all proofs to App. A.
@@
satisfies Assn. 2 with $L_{0}=0$, $L_{1}=8\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$, and Assn. 3 with $\nu=8\max_{i}\left\|x_{i}\right\|$, $\omega=0$.
Note that the logistic regression objective is also uniform smooth, meaning that it simultaneously satisfies Assn. 2 with $L_{0}=\frac{1}{4n}\,\lambda_{\max}[X^{T}X]$ and $L_{1}=0$, where $X\in\mathbb{R}^{n\times d}$ is the corresponding feature matrix. On the other hand, binary classification with the exponential loss is not uniform smooth on an unbounded domain, but satisfies Assn. 2 with $L_{0}=0$, $L_{1}=8\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$ (see Prop. 6 in App. A).
Next, we show that the non-convex objective corresponding to the generalized linear model (GLM) with a logistic link function also satisfies Assn. 1, 2 and 3 . In particular, we prove the following proposition in App. A.
@@
Consider $n$ points where $x_{i}\in\mathbb{R}^{d}$ are the features and $y_{i}\in[0,1]$ are the corresponding labels. If $\pi_{i}(\theta)=\sigma(\langle x_{i},\theta\rangle):=\frac{1}{1+\exp(-\langle x_{i},\theta\rangle)}$, the GLM objective,
$f(\theta)=\frac{1}{2n}\sum_{i=1}^{n}\left(\pi_{i}(\theta)-y_{i}\right)^{2}\,,$ (4)
satisfies Assn. 2 with $L_{0}=\frac{9}{16}\,\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$, $L_{1}=9\,\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$ and Assn. 3 with $\nu=9\max_{i}\left\|x_{i}\right\|$, $\omega=\max_{i}\left\|x_{i}\right\|$.
Finally, in App. A, we show that the objective for softmax policy gradient (Mei et al., 2020) also satisfies the required assumptions for multi-armed bandits and tabular Markov decision processes, and we study these objectives in Sec. 5.
Connection to non-uniform smoothness in Zhang et al. (2019): Assn. 2 and 3 are related to the non-uniform smoothness condition proposed in Zhang et al. (2019) and analyzed in Zhang et al. (2020); Li et al. (2023); Chen et al. (2023); Vankov et al. (2024); Gorbunov et al. (2024). In particular, we prove the following result in App. A.
Proposition 3.
For a non-negative, twice-differentiable function $f$, if $f$ is $(L_{c},L_{g})$ non-uniform smooth according to Zhang et al. (2019) i.e.
$\left\|\nabla^{2}f(\theta)\right\|\leq L_{c}+L_{g}\,\left\|\nabla f(\theta)\right\|\,,$
then, it is satisfies Assn. 2 with $L_{0}=L_{c}+L_{g}\,\sqrt{2L_{c}}$, and $L_{1}=L_{g}\,\left(8L_{g}+\sqrt{2L_{c}}\right)$ and Assn. 3 with $\nu=8L_{g}+\sqrt{2L_{c}}$ and $\omega=\sqrt{2L_{c}}$.
Hence, the $(L_{c},L_{g})$ non-uniform smoothness in Zhang et al. (2019) implies Assn. 2 and 3. Consequently, our subsequent results also apply to non-negative, twice-differentiable, $(L_{c},L_{g})$ non-uniform smooth functions.
In order to further motivate Assn. 2 and 3, consider the logistic regression example in Prop. 1. For logistic regression, the loss corresponding to a single point satisfies the notion of $(L_{c},L_{g})$ non-uniform smoothness with $L_{c}=0$ and $L_{g}=\left\|x_{i}\right\|$ (Gorbunov et al., 2024, Example 1.6). However, the finite-sum over $n$ points ($f(\theta)$ in Eq. 3) does not necessarily satisfy this assumption with $L_{c}=0$ (see Prop. 8 for a simple example with $n=2$). Consequently, unlike the result in Prop. 1, using the $(L_{c},L_{g})$ condition in conjunction with Prop. 3 does not directly imply that $L_{0}=0$. In Sec. 4, we will see that, for logistic regression, having $L_{0}=0$ is the key to achieving fast convergence for GD-LS. This further justifies our alternate definition of non-uniform smoothness.
Now that we have motivated Assn. 1, 2 and 3, in the next section, we consider minimizing such non-uniform smooth functions using gradient descent with Armijo line-search.